7) A product is made of four identical components. In order for the product to function properly all of the components must function. All components have equal performance probabilities. The product probability of performance is 0.98. What is the minimum probability of performance by each of the four components?

a)         0.9       b)         0.854   c)         0.879   d)         0.638   e)         0.995
Answer #1

let minimum probability of performance by each of the four components=a

P(system work) =P(all 4 works )=(a)4 =0.98

a =0.981/4 =0.995

option E